OpenStudy (anonymous):

Consider the real-valued function given by the following rule. What is the largest domain on which this rule makes sense?

OpenStudy (anonymous):

\[f(x) = \sqrt[4]{4-x^4}\]

OpenStudy (danjs):

The term under the root must be greater than or equal to zero

OpenStudy (danjs):

\[4-x^4 = -(x^2 +2)(x^2-2)\]

OpenStudy (danjs):

Or you can just say \[4 - x^4 \ge0 ~~~so~~~~x^4 \le 4\] \[x \le 4^{1/4}\] \[x \le \sqrt{2}\]

OpenStudy (danjs):

+ or - square root of 2 so x has to be in between + and - square root 2

OpenStudy (danjs):

\[-\sqrt{2}\le x \le \sqrt{2}\]